MODS1 Commissioning Run 6: UTC 2011 March 17

Observer(s): R. Pogge (OSU) and P. Garnavich (ND)
Telescope Operator: S. Allanson (LBTO)
AO Support: A. Rakich, J. Hill, & D. Miller Remote from Tucson
Telescope Support: J. Urban (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

Observed all night, through thin cirrus and bright moonlight. The first major portion of the night was devoted to active optics work. Andrew collected a lot of data and we made a lot of progress. Looks like the major problems we had are getting solved. Started some of the final verification tests while waiting for the telescope to settle enough for Andrew to start. More are on the schedule for subsequent nights.

The last third of the night we reached another milestone: first multi-object spectroscopy using a test mask of the Abell 1689 cluster - we've finally put the MO in MODS. This was also our first attempt to align a 2D slit mask, and there was much bumping around as we tried to get our alignment code running. It was a good start and we gathered a lot of experience and data that will help us on the next steps on following nights.

No opportunity yet to exercise the new target-acquisition scripts, but we did all of our end-of-night calibrations using scripts developed on the last run. Looks like our estimates of exposure times and filter combinations for lamps and flats are robust.

Preparations

Daytime work on MODS1 on March 14 and 15. A quick summary
  • Replaced the damaged bottom cover with a new cover that allows better access to the collimator mirror covers
  • Restored guide camera function (including a full hardware check and software bug fix)
  • Realignment of the WFS camera, but only modest improvement on vignetting
  • Verified that the TEC on the guide camera is not functioning, will replace it with the MODS2 guide camera head in April after the latter has glycol fittings installed and pressure/leak tested.
  • Tested MODS last night with BP10 and verified basic function including on- and off-axis active presets

We are ready to resume commissioning, with the first night devoted to remaining AGw issues (updates to the guide star acquisition and WFS analysis modules).

Marco is not coming up until Thursday.

Details

All times are UTC unless otherwise noted.

2330 - Warm on the mountain today, 8-10°C, light wind 5-8m/s from the SW, humidity ~35%. Some cirrus around, but otherwise a nice day. There are some cloud bands way off to the west that are heading our way, though...

0140 - Sunset, we've been open for a while with vent doors to get equilibrated.

0200 - Setting to BS9126 for on-axis WFS setup. Still bright out, and swimmy, so we'll sit here for a while and see what we can see. Andrew is on from Tucson to supervise this part of the evening.

Seeing is swimmy, maybe 1.5-2.0 arcsec and variable. Looking at the sky the stars are a little twinkly at the meridian. Cool view of Jupiter and Mercury over on the western horzion.

0245 - Taking over from Andrew. Doing the along-slit offset. Seeing is 1.5-2 arcsec, so not the greatest.

Running the slitscan sequence on BS9126 using a 4-arcmin offset guide star and the fat 5-arcsec slit. first sequence was 0/+30/+15/-15/-30 (detxy), but seeing stank.

Second sequence seeing was improving, now doing 0/+20/+10/-10/-20 to stay inside the slit vertically (oops), and a slightly shorter exposure.

During these data we saw a recurrance of the Y-direction "fox tails" distortion of the images. This is what we were getting back in previous months when the WFS analysis code still had an incorrect coordinate transformation. Andrew is checking with the software folks to find out what version we're really running. Something else going on?

0347 - Back to on-axis, checking collimation.

We're still seeing aberrated PSFs. The problem might be related to the WFS camera realignment from earlier this week and the current S-H algorithm for finding the Shack spots may be having problems finding the spots on different parts of the GRIN (having aligned it to where Doug & Andrew wanted it on-axis).

Trying to reach Doug...

0400 Doug working on the problem, will check soon. Also fixed a problem with timeout while waiting for .fits and .info files, will see if theproblem is fixed soon...

0412 Updated WFShotspot_X to 382.5 and WFShotspot_Y to 252.0 in OSU_L.cfg as just measured by Doug on a MODS on-axis pupil.

0425 running field_collect on BS9126

0450 mirror went down, aborted IDL on GS 10 (# 5 in sequence). Will restart on START=5 when mirror comes back up...

0525 major TCS problems, still waiting...

0605 have been underway for 15 minutes or so. Star # 23 (#7 in list) for BS9126 is out of range of the stage, restarting at # 8 on the list.

0610 AGAIN lost connection to mountain and the only recovery seems to be to log off and log back on, 15 -20 minutes downtime.

0625 star 24 also out of range. Doug we need star lists for MODS with different acceptable ranges than Lucifer. Radius should be no more 4.5 arc minutes max.

0703 Going to Stone I for 6 more stars for field collect, then will do field check on this field as well. Seeing consistant though most of tonight around 1"-1.5".

0720 A probable explanation for the consistant positive focus correction for off-axis measurements is that at the moment the MODS agw is not correctly taking into account the offset of the WFS from stage position, will look at the data in more detail tomorrow, looking for a smooth function of focus error with off-axis distance.

0735 finished field collect, starting field_check on STONE I.

0745 our first PA rotation off the y axis for GS8 (star 1 in the sequence) returned the following error:

# Y-Axis Position Angle: 353.59

request = irc PresetTelescope 413.59 POSITION ACTIVE left; echo $? ===> Error in irc_presetTelescope err = Resmsg[0] = PresetTelescope execution failed: PresetTelescope: Parameter 'rotationAngle' out of range, Possible values are: [-2PI, 2PI] radians % Program caused arithmetic error: Floating illegal operand

Looks like if the program had requested a PA of 53 degrees Trying star 2 on the list.

0800 failed this time because x value too big with 60 degree pa offset.

0813 Doug fixes code to a) limit pa offsets to +/- 40 and b) make requested pa modulo 2Pi.

0826 Doug fixes code again to make a) from above work properly. Restarting field_check.

0845 running smoothly now, IDL and GCS outputs seem to be in good agreement too.

0930 Star five was out of X range by 4.5 mm, going on to star 6.

0945 That's that for now, over to Rick.

0952 Acquiring the Abell 1689 field for the first MOS test mask.

The modsalign.py script is giving some odd errors, resorting to at least an xy shift by hand with XVista. Got reasonable alignment at middle third or so of the field, but can see residual rotation of <1° in the corners.

1024 - Got the first MODS multi-object spectrum, 300s blue+red in the Abell 1689 field!

1100 - Messing about with mask alignment on the Abell 1689 field. Some funny python script problems getting in our way. Paul should be up soon so we'll ask him what's up.

1200 - Interesting lesson learned - don't mail python scripts as attachments. You really don't want to know how truly weird it was.

Anyway, even with a pure translation and no rotation we got pretty good alignment. We took a number of acquisition images with and without the slit in the r, i, and z filters. This will give us lots of data to use to exercise modsalign.py and shake the bugs out. Seeing was variable, 0.8-1.2 arcsec, worse as we descended to lower airmass.

1230 - A minor mistake at the end, we took slitless spectra with the grating. Oops. We were getting pretty tired towards the end there. Running calibrations at the end of the night to give us a complete look at an MOS mask.

Closing up and putting MODS and ourselves to bed soon.

-- RichardPogge - 16 Mar 2011
Topic revision: r9 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback